# Double-coset zeta functions for groups acting on trees

### v1

@@ 4.1. Cosets and geodesics
$\begin{split}\varphi_{v,t}(gG_{t})&:=\left\{\begin{array}[]{cc}[v,g\cdot t],&\text{if }[v,g\cdot t]\text{ exists in }T;\\ {[v,g\cdot\bar{t}]},&\text{if }[v,g\cdot\bar{t}]\text{ exists in }T;\end{array}\right.\quad\text{and}\\ \varphi_{e,t}(gG_{t})&:=\left\{\begin{array}[]{cc}{[e,g\cdot t]},&\text{if }g\cdot t\in T_{\geq e};\\ {[e,g\cdot\bar{t}]},&\text{if }g\cdot\bar{t}\in T_{\geq e};\\ {[\bar{e},g\cdot t]},&\text{if }g\cdot t\in T_{\geq\bar{e}};\\ {[\bar{e},g\cdot\bar{t}]},&\text{if }g\cdot\bar{t}\in T_{\geq\bar{e}.}\end{array}\right.\end{split}$
@@
For every $g\in G$, we observe that
@@ 4.2. The case of weakly locally $\infty$-transitive actions on trees
Let $(G,T)$ be a weakly locally $\infty$-transitive group action on a tree. Denote by $\pi\colon T\to\Gamma=G\backslash T$ the quotient map of $(G,T)$, and let $\omega$ be the standard edge weight on $\Gamma$. Given $u_{1},u_{2}\in\Gamma$ and $t_{1}\in\pi^{-1}(u_{1})$, let $\mathcal{P}^{\mathrm{lift}}_{\Gamma,t_{1}}(u_{1}\to u_{2})$ be the set of all paths $\mathfrak{p}\in\mathcal{P}_{\Gamma}(u_{1}\to u_{2})$ that can be lifted via $\pi$ to a geodesic in $T$ from $t_{1}$.
@@
In the hypotheses before, let $v\in VT$ with $\pi(v)=c$ and $e\in ET$ with $\pi(e)=a$. According to Fact 4.1, we have the following bijections for every $t\in VT$ with $\pi(t)=u$:
@@
Here the sets $A$ and $U$ are according to Notation 2.1.
@@
Clearly, $\pi_{v,t}$ is $G$-equivariant and surjective. Moreover, since the $G$-action on $T$ is weakly locally $\infty$-transitive, for every $g\in G$ the $\pi_{v,t}$-fibre of $\pi([v,g\cdot t])$ is $G_{v}\cdot[v,g\cdot t].$ Thus $\pi_{v,t}$ induces a $1$–to–$1$ map
@@
In view of Proposition 4.6, we introduce a weight on the paths of $G\backslash T$ which extends the standard edge weight $\omega$ and such that $|G_{t_{1}}gG_{t_{2}}/G_{t_{2}}|$ coincides with the weight of $\Psi_{t_{1},t_{2}}(G_{t_{1}}gG_{t_{2}})$, for all $t_{1},t_{2}\in T$. Such a weight can be defined even in the following more general framework.
@@
Let $\Gamma$ be a graph with a function $\omega\colon E\Gamma\to\mathbb{Z}_{\geq 1}$ (called edge weight). Define two functions $N_{\mathrm{edg}}=N_{\mathrm{edg}}^{\omega},N_{\mathrm{vert}}=N_{\mathrm{vert}}^{\omega}\colon\mathcal{P}_{\Gamma}\to\mathbb{Z}_{\geq 0}$ as follows. For every path $\mathfrak{p}$ in $\Gamma$, let
@@
Let $\Gamma$ be the quotient graph of a group action on a tree $(G,T)$, and denote by $\omega\colon E\Gamma\to\mathbb{Z}_{\geq 1}$ its standard edge weight. For every path $\mathfrak{p}=(a_{1},\ldots,a_{n})$ in $\Gamma$ of positive length and for every $e\in ET$ with $\pi(e)=a_{1}$, one checks that
@@
These observations apply if in particular $\Gamma$ is the underlying graph of a local action diagram $\Delta=(\Gamma,(X_{a}),(G(c)))$, and if $T=T(\Delta,\iota,c_{0})$ is a standard $\Delta$-tree (cf. Section 3.4.1). In this case, for path $\mathfrak{p}=(a_{1},\ldots,a_{n})$ in $\Gamma$ of positive length and every $e\in ET^{+}$ with $\pi(e)=a_{1}$, we have
@@
(cf. Lemma 3.5). Thus a reduced path $\xi$ in $(\Delta,\iota)$ can be lifted to a geodesic in $T$ if, and only if, $N_{\mathrm{edg}}^{\omega}(\mathfrak{p}_{\xi})\geq 1$.
@@
Under the hypotheses of Proposition 4.3, suppose that $\omega(a)\geq 2$ for every $a\in E\Gamma$. Then every path $\mathfrak{p}$ can be lifted to a geodesic in $T$ via $\pi$ (cf. Remark 4.5.1). In particular, the bijections in Proposition 4.3 are onto $\mathcal{P}_{\Gamma}(c\to U)$ and $\mathcal{P}_{\Gamma}(A\to U)$, respectively.
@@
Let $(G,T)$ be a weakly locally $\infty$-transitive group action on a tree with quotient graph $\Gamma$. Assume that the standard edge weight $\omega$ on $\Gamma$ takes finite values. Then, for all $g\in G$ and $t_{1},t_{2}\in T$, we have
$|G_{t_{1}}gG_{t_{2}}/G_{t_{2}}|=|G_{t_{1}}:G_{\varphi_{t_{1},t_{2}}(gG_{t_{2}})}|=\left\{\begin{array}[]{ll}N_{\mathrm{vert}}^{\omega}(\Psi_{t_{1},t_{2}}(G_{t_{1}}gG_{t_{2}})),&\text{if }t_{1}\in VT;\\ N_{\mathrm{edg}}^{\omega}(\Psi_{t_{1},t_{2}}(G_{t_{1}}gG_{t_{2}})),&\text{if }t_{1}\in ET.\end{array}\right.$
@@ 4.3. The case of (P)-closed actions on trees
Let $\Delta$ be a local action diagram and $(T=T(\Delta,\iota,c_{0}),\pi,\mathcal{L})$ be the standard $\Delta$-tree associated to an inversion map $\iota$ in $\Delta$ and a chosen $c_{0}\in V\Gamma$ (cf. Section 3.4.1). Let $G\leq U(\Delta,\iota,c_{0})$ be a subgroup acting on $T$ with local action diagram $\Delta$. For the definition of $U(\Delta,\iota,c_{0})$, see Section 3.4.2.
@@
Let $G\leq U(\Delta,\iota,c_{0})$ and $T=T(\Delta,\iota,c_{0})$ as before, and denote by $v_{0}$ the root of $T$. Consider also $e\in o^{-1}(v_{0})$ with $\mathcal{L}(e)=x$ and $t\in T$ with $\pi(t)=u$. Let $\varphi_{u,t},\varphi_{e,t}$ be the maps introduced in Fact 4.1, and denote by $\mathcal{L}$ the map defined in (3.5). Then the following two maps are bijective:
@@
For $x,y\in X$ note that $\mathcal{W}(x,y)\neq 0$ if, and only if, $(x,y)$ is a path in $\Delta$. More generally, given a sequence $\xi$ of elements of $X$ we have $\mathcal{W}(\xi)\neq 0$ if, and only if, $\xi$ is a path in $\Delta$.
@@
Assume the hypotheses of the section, set $G=U(\Delta,\iota,c_{0})$ and denote by $\omega$ and $\mathcal{W}$ be the standard edge weights on $\Gamma$ and $\Delta$, respectively. Hence, for every $e\in ET$ we have
@@
Let $\Delta=(\Gamma,(X_{a}),(G(c)))$ be a local action diagram and consider $a,b\in E\Gamma$ with $t(a)=c=o(b)$. Assume that, for every $x\in X_{a}$, the group ${G(c)}_{\iota(x)}$ acts transitively on $X_{b}\setminus\{\iota(x)\}$. Then, for all $x\in X_{a}$ and $y\in X_{b}\setminus\{\iota(x)\}$ we have
@@
By Proposition 3.1, for every $11\leq i\leq n-$ 1 we have
@@
Let $G=U(\Delta,\iota,c_{0})$ and denote by $v_{0}$ be the root of $T=T(\Delta,\iota,c_{0})$. Let $e\in ET\setminus ET^{+}$ with $t(e)=v_{0}$, and consider $t\in T$ such that $[v_{0},t]$ is defined. Set $\mathfrak{p}_{t}:=e\cdot[v_{0},t]=(e_{1}=e,e_{2},\ldots,e_{n})$, $\mathcal{L}(\bar{e})=x_{1}$ and $\mathcal{L}(\mathfrak{p}_{t})=(\iota(x_{1}),x_{2},\ldots,x_{n})$. Then
@@
From Remark 3.4.3 observe that $e_{2},\ldots,e_{n}\in ET^{+}$. Hence, Proposition 3.1 and Remark 4.9.1 yield
@@
If $e_{1}\in ET^{+}$, by Remark 3.4.3 we may take $g=1$ and Proposition 4.11 applies. Assume now that $e_{1}\in ET\setminus ET^{+}$ and let $(f_{1},\ldots,f_{r}=\bar{e}_{1})$ be the geodesic from the root $v_{0}$ of $T$ to $\bar{e}_{1}$. Set $\mathcal{L}(f_{i})=x_{i}$ for every $1\leq i\leq r$. In particular, $\mathcal{L}(\bar{e}_{1})=x_{r}$ and then $\mathcal{L}(e_{1})=\iota(x_{r})$ (cf. Remark 3.4.1). Since $|X_{\pi(e_{1})}|\geq 2$, there is $y\in X_{\pi(e_{1})}\setminus\{\iota(x_{r})\}$ such that $(x_{1},\ldots,x_{r},y)$ is a reduced path in $(\Delta,\iota)$. Then there is $f\in ET^{+}$ such that $o(f)=o(e_{1})$ and $t(f)$ corresponds to $(x_{1},\ldots,x_{r},y)$. Since $\mathcal{L}(f)=y,\mathcal{L}(e_{1})\in X_{\pi(e_{1})}$, there is $g\in G_{o(e_{1})}$ such that $f=g\cdot e_{1}$. By Remark 3.4.3, every edge of $g\cdot\mathfrak{p}=(g\cdot e_{1},\ldots,g\cdot e_{n})$ belongs to $ET^{+}$. Moreover,
@@ 5. Double-coset zeta functions for groups acting on trees
This section deals with the convergence and explicit formulae for the relevant double-coset zeta functions for groups acting on trees. We introduce a family of properties (labelled with positive integers) on group actions on trees (cf. Section 5.2). In Proposition 5.5, we exploit that one of these properties is satisfied to deduce that the group has polynomial double-coset growth with respect to vertex or edge stabilisers. The latter result can be refined to a characterisation in case that the action is weakly locally $\infty$-transitive (cf. Theorem 5.6) or (P)-closed (cf. Theorem 5.7). In these two cases, we also provide explicit formulae for the relevant double-coset zeta functions in terms of the local data introduced in Section 4 (cf. Theorems 5.12 and 5.19).

### v2

@@ 4.1. Cosets and geodesics
[…]
@@
For every $g\in G$ we observe that
@@ 4.2. The case of weakly locally $\infty$-transitive actions on trees
Let $(G,T)$ be a weakly locally $\infty$-transitive group action on a tree. Denote by $\T\pi\colon T\to\Gamma=G\backslash$ T the quotient map of $(G,T)$, and let $\omega$ be the standard edge weight on $\Gamma$. Given $u_{1},u_{2}\in\Gamma$ and $t_{1}\in\pi^{-1}(u_{1})$, let $\mathcal{P}^{\mathrm{lift}}_{\Gamma,t_{1}}(u_{1}\to u_{2})$ be the set of all paths $\mathfrak{p}\in\mathcal{P}_{\Gamma}(u_{1}\to u_{2})$ that can be lifted to a geodesic in $T$ from $t_{1}$ via $\pi$.
@@
In the hypotheses before, let $v\in VT$ with $\pi(v)=c$ and $e\in ET$ with $\pi(e)=a$. According to Fact 4.1, we have the following bijections for every $t\in T$ with $\pi(t)=u$:
@@
Here the sets $A$ and $U$ are as in Notation 2.1.
For all $t_{1},t_{2}\in T$ and all $h\in G_{t_{1}}$, note that
$\pi(\varphi_{t_{1},t_{2}}(hgG_{t_{2}}))=\pi(h\cdot\varphi_{t_{1},t_{2}}(gG_{t_{2}}))=\pi(\varphi_{t_{1},t_{2}}(gG_{t_{2}}))$
and thus the map $\Psi_{t_{1},t_{2}}$ is well-defined.
@@
Clearly, $\pi_{v,t}$ is $G$-equivariant and surjective. Moreover, since the $G$-action on $T$ is weakly locally $\infty$-transitive, for every $g\in G$ the $\pi_{v,t}$-fibre of $\pi([v,g\cdot t])$ is $G_{v}\cdot[v,g\cdot t].$ Thus $\pi_{v,t}$ induces a bijective map
@@
[…]
@@
Let $\Gamma$ be a graph with a function $\omega\colon E\Gamma\to\mathbb{Z}_{\geq 1}$ called edge weight. Define two functions $N_{\mathrm{edg}}=N_{\mathrm{edg}}^{\omega},N_{\mathrm{vert}}=N_{\mathrm{vert}}^{\omega}\colon\mathcal{P}_{\Gamma}\to\mathbb{Z}_{\geq 0}$ as follows. For every path $\mathfrak{p}$ in $\Gamma$, let
@@
Let $\Gamma$ be the quotient graph of a group action on a tree $(G,T)$, and denote by $\omega\colon E\Gamma\to\mathbb{Z}_{\geq 1}$ its standard edge weight.
For every path $\mathfrak{p}=(a_{1},\ldots,a_{n})$ in $\Gamma$ of positive length and for every $e\in ET$ with $\pi(e)=a_{1}$, one checks that
@@
These observations apply if in particular $\Gamma$ is the underlying graph of a local action diagram $\Delta=(\Gamma,(X_{a}),(G(c)))$, and $T=T(\Delta,\iota,c_{0})$ is a standard $\Delta$-tree (cf. Section 3.4.1). In this case, given a path $\mathfrak{p}=(a_{1},\ldots,a_{n})$ in $\Gamma$ of positive length and for every $e\in ET^{+}$ with $\pi(e)=a_{1}$, we have
@@
(cf. Lemma 3.5). Thus, a reduced path $\xi$ in $(\Delta,\iota)$ of positive length can be lifted to a geodesic in $T$ if, and only if, $N_{\mathrm{edg}}^{\omega}(\mathfrak{p}_{\xi})\geq 1$.
@@
Assume the notation of Remark 4.5.1 and suppose that $\omega(a)\geq 2$ for every $a\in E\Gamma$. Then every path $\mathfrak{p}$ can be lifted to a geodesic in $T$ via $\pi$ (cf. Remark 4.5.1). In particular, the bijections in Proposition 4.3 are onto $\mathcal{P}_{\Gamma}(c\to U)$ and $\mathcal{P}_{\Gamma}(A\to U)$, respectively.
@@
Let $(G,T)$ be a weakly locally $\infty$-transitive group action on a tree with quotient graph $\Gamma$. Assume that the standard edge weight $\omega$ on $\Gamma$ takes finite values. Then, for all $g\in G$ and $t_{1},t_{2}\in T$,
$|G_{t_{1}}gG_{t_{2}}/G_{t_{2}}|=|G_{t_{1}}:G_{\varphi_{t_{1},t_{2}}(gG_{t_{2}})}|=\left\{\begin{array}[]{ll}N_{\mathrm{vert}}^{\omega}(\Psi_{t_{1},t_{2}}(G_{t_{1}}gG_{t_{2}})),&\hskip-5.69046pt\text{if }t_{1}\in VT;\\ N_{\mathrm{edg}}^{\omega}(\Psi_{t_{1},t_{2}}(G_{t_{1}}gG_{t_{2}})),&\hskip-5.69046pt\text{if }t_{1}\in ET.\end{array}\right$.
@@ 4.3. The case of (P)-closed actions on trees
Let $\Delta$ be a local action diagram and let $(T=T(\Delta,\iota,c_{0}),\pi,\mathcal{L})$ be the standard $\Delta$-tree associated to an inversion map $\iota$ in $\Delta$ and a chosen $c_{0}\in V\Gamma$ (cf. Section 3.4.1). Let $G\leq U(\Delta,\iota,c_{0})$ be a subgroup acting on $T$ with local action diagram $\Delta$. For the definition of $U(\Delta,\iota,c_{0})$, see Section 3.4.2.
@@
Let $G\leq U(\Delta,\iota,c_{0})$ and $T=T(\Delta,\iota,c_{0})$ be as before, and denote by $v_{0}$ the root of $T$. Consider also $e\in o^{-1}(v_{0})$ with $\mathcal{L}(e)=x$ and $t\in T$ with $\pi(t)=u$. Let $\varphi_{u,t},\varphi_{e,t}$ be the maps introduced in Fact 4.1, and denote by $\mathcal{L}$ the map defined in (3.5). Then the following two maps are bijective:
@@
For $x,y\in X$ note that $\mathcal{W}(x,y)\neq 0$ if, and only if, $(x,y)$ is a path in $\Delta$. More generally, given a sequence $\xi$ of elements of $X$, we have $\mathcal{W}(\xi)\neq 0$ if and only if $\xi$ is a path in $\Delta$.
@@
Assume the hypotheses of the section, let $G=U(\Delta,\iota,c_{0})$ and denote by $\omega$ and $\mathcal{W}$ the standard edge weights on $\Gamma$ and $\Delta$, respectively. For every $e\in ET$ we have
@@
Let $\Delta=(\Gamma,(X_{a}),(G(c)))$ be a local action diagram and consider $a,b\in E\Gamma$ with $t(a)=c=o(b)$. Assume that, for every $x\in X_{a}$, the group ${G(c)}_{\iota(x)}$ acts transitively on $X_{b}\setminus\{\iota(x)\}$. Note that $X_{b}\setminus\{\iota(x)\}=X_{b}$ unless $b=\bar{a}$. Then, for all $x\in X_{a}$ and $y\in X_{b}\setminus\{\iota(x)\}$, we have
@@
By Lemma 3.1, for every $1\leq i\leq n-1$ we have
@@
Let $G=U(\Delta,\iota,c_{0})$ and denote by $v_{0}$ the root of $T=T(\Delta,\iota,c_{0})$. Let $e\in ET\setminus ET^{+}$ with $t(e)=v_{0}$, and consider $t\in T$ such that $v_{0},t]$ is defined. Set $\mathfrak{p}_{t}:=e\cdot[v_{0},t]=(e_{1}=e,e_{2},\ldots,e_{n})$, $\mathcal{L}(\bar{e})=x_{1}$ and $\mathcal{L}(\mathfrak{p}_{t})=(\iota(x_{1}),x_{2},\ldots,x_{n})$. Then
@@
From Remark 3.4.3 observe that $e_{2},\ldots,e_{n}\in ET^{+}$. Hence, Lemma 3.1 and Remark 4.9.1 yield
@@
If $e_{1}\in ET^{+}$, by Remark 3.4.3 we may take $g=1$ and Proposition 4.11 applies. Assume now that $e_{1}\in ET\setminus ET^{+}$ and let $(f_{1},\ldots,f_{r}=\bar{e}_{1})$ be the geodesic from the root $v_{0}$ of $T$ to $\bar{e}_{1}\in ET^{+}$. Set $\mathcal{L}(f_{i})=x_{i}$ for every $1\leq i\leq r$. In particular, $\mathcal{L}(\bar{e}_{1})=x_{r}$ and then $\mathcal{L}(e_{1})=\iota(x_{r})$ (cf. Remark 3.4.1). Since $|X_{\pi(e_{1})}|\geq 2$, there is $y\in X_{\pi(e_{1})}\setminus\{\iota(x_{r})\}$ such that $(x_{1},\ldots,x_{r},y)$ is a reduced path in $(\Delta,\iota)$. Then there is $f\in ET^{+}$ such that $o(f)=o(e_{1})=t(\bar{e}_{1})$, and $t(f)$ corresponds to $(x_{1},\ldots,x_{r},y)$ (cf. (3.4)). Since both $\mathcal{L}(f)=y$ and $\mathcal{L}(e_{1})$ belong to $X_{\pi(e_{1})}$, there is $g\in G_{o(e_{1})}$ such that $f=g\cdot e_{1}$. By Remark 3.4.3, every edge of $g\cdot\mathfrak{p}=(g\cdot e_{1},\ldots,g\cdot e_{n})$ belongs to $ET^{+}$. Moreover,
@@ 5. Double-coset zeta functions for groups acting on trees
This section deals with the convergence and explicit formulae for the relevant double-coset zeta functions for groups acting on trees. We introduce a family of properties (labelled with positive integers) on group actions on trees (cf. Section 5.2). In Proposition 5.7, we exploit that one of these properties is satisfied to deduce that the group has polynomial double-coset growth with respect to vertex or edge stabilisers. The latter result can be refined to a characterisation in case that the action is weakly locally $\infty$-transitive (cf. Theorem 5.8) or (P)-closed (cf. Theorem 5.9). In these two cases, we also provide explicit formulae for the relevant double-coset zeta functions in terms of the local data introduced in Section 4 (cf. Theorems 5.14 and 5.21).
We remind that we are following Convention 3.0.1 and Convention 3.1.1.
@@ 5.1. From double-cosets to cosets
Remark 5.0.1.
Let $\Delta$ be a local action diagram and $\mathbb{T}=(T,\pi,\mathcal{L})$ be $\Delta$-tree. Consider an inversion $\iota$ in $\Delta$ and denote by $(T(\Delta,\iota,c_{0}),\pi_{0},\mathcal{L}_{0})$ the standard $\Delta$-tree associated to $\iota$ and some $c_{0}\in V\Gamma$. By Theorem 3.8(i), there is a graph isomorphism $\phi\colon T\to T(\Delta,\iota,c_{0})$ such that $\pi=\pi_{0}\circ\phi$ and $U(\Delta,\iota,c_{0})=\phi U(\Delta,\mathbb{T})\phi^{-1}$. Set $G=U(\Delta,\mathbb{T})$ and $H=U(\Delta,\iota,c_{0})$. Then, for all $t\in T$, the following map is bijective:
$|G_{t_{1}}:G_{t_{1}}\cap gG_{t_{2}}g^{-1}|=|H_{\phi(t_{1})}:H_{\phi(t_{1})}\cap hH_{\phi(t_{2})}h^{-1}|.$
Hence,
$a_{n}(G,G_{t_{1}},G_{t_{2}})=a_{n}(H,H_{\phi(t_{1})},H_{\phi(t_{2})}),\quad\forall\,n\geq 1$
as well as
$\zeta_{G,G_{t_{1}},G_{t_{2}}}(s)=\zeta_{H,H_{\phi(t_{1})},H_{\phi(t_{2})}}(s)$
when the series before are defined.
$\begin{split}a_{n}(G,G_{v},G_{t})&=a_{n}(H,H_{v_{0}},H_{\phi(t)}),\quad\forall\,n\geq 1;\quad\text{and}\\ \zeta_{G,G_{v},G_{t}}(s)&=\zeta_{H,H_{v_{0}},H_{\phi(t)}}(s).\end{split}$
Moreover, for all $e\in o^{-1}(v)$, we have $\phi(e)\in o^{-1}(v_{0})$ and then, for all $t\in T$,
$\begin{split}a_{n}(G,G_{e},G_{t})&=a_{n}(H,H_{\phi(e)},H_{\phi(t)}),\quad\forall\,n\geq 1;\quad\text{and}\\ \zeta_{G,G_{e},G_{t}}(s)&=\zeta_{H,H_{\phi(e)},H_{\phi(t)}}(s).\end{split}$